a)Calculate the wavelength of light in a vacuum that has a frequency of 8.5 x 109 Hz.
5.45 m

(b) What is its wavelength in glycerine? (The index of refraction of glycerine is 1.473.)
3.70 × 10 -9 m

(c) Calculate the energy of one photon of such light in vacuum. Express the answer in electron volts
228 eV
1
Expert's answer
2018-10-05T11:23:29-0400
a) Calculate the wavelength of light in a vacuum that has a frequency of 8.5 x 109 Hz.
Solution
λ=c/f=(3∙〖10〗^8)/(8.5∙〖10〗^9 )=0.0353 m=3.53 cm.

(b) What is its wavelength in glycerin? (The index of refraction of glycerin is 1.473.)
Solution
λ_g=1/n c/f=1/1.473 (3∙〖10〗^8)/(8.5∙〖10〗^9 )=0.0240 m=2.40 cm.

(c) Calculate the energy of one photon of such light in vacuum. Express the answer in electron volts
Solution
E=hc/λ=(1240 eVnm)/(0.0353∙〖10〗^9 nm)=3.51∙〖10〗^(-5) eV.